OpenStudy (anonymous):

Help! Find the area of the region bounded by y=x^3-4x^2+3x and the x-axis

OpenStudy (amistre64):

x(x^2 -4x +3) x(x-3)(x-1) x = 0,1,3

OpenStudy (amistre64):

your gonna have a negative area and a positive area.... they either want it as + and - or as a total absolute value...

OpenStudy (amistre64):

x^4/4 -4x^3/3 +3x^2/2 ; [0,3] at x = 0 we get 0 soo thats redundant at 3 we get 3^4/4 -4(3)^3/3 +3(3)^2/2

OpenStudy (amistre64):

-9 if they want a +,- area value....

OpenStudy (amistre64):

-9/4.... that is

OpenStudy (amistre64):

the othe way is: 5/12 + 32/12 = 37/12...i think

OpenStudy (amistre64):

a + b = -9/4 a = 5/12 -b = -9/4 - 5/12 b = 27+5/12 b = 32/12 a+b = 5+32/12 = 37/12.....

OpenStudy (amistre64):

a + b = -9/4 a = 5/12 b = -9/4 - 5/12 b = -27-5/12 b = |-32/12| a+b = 5+32/12 = 37/12 ....same results :)

OpenStudy (anonymous):

ok that makes a lot more sense to me! thanks amistre